# Fourier dimension at $C^{1+\alpha}$ regularity

### v2

@@ 1.2. Results in $C^{1+\alpha}$ regularity
If the $C^{1+\alpha}$ IFS has finitely many branches, then (MNL) implies (QNL).
Thus by Theorem 1.1, MNL suffices to obtain power Fourier decay for the conformal measure $\mu$ on $F_{\Phi}$. We now construct $C^{1+\alpha}$ IFSs satisfying MNL, yielding fractal equilibrium states with power Fourier decay. As far as we know, these are the first examples of $C^{1+\alpha}$ IFSs not conjugated to linear IFSs admitting power Fourier decay.
The construction proceeds via fixed point theory in the moduli space of IFSs parametrized by equilibrium states, ensuring the required non-concentration to implement the proof of Theorem 1.1. The key input is the existence of a branch whose log-derivative has $\alpha$–Hölder lower regularity, as furnished by a Cantor-type function with jumps precisely on the attractor.
Theorem 1.4.
which depend on the equilibrium state $\mu$, to produce measures with the Measure Non-Linearity (MNL) property; Theorem 1.1 then applies. Let us now discuss on the MNL condition and prospects of the theory.
Remark.
  • (1)
    MNL property and autosimilar temporal distance function. The Cantor-type structure of $\tau$ in our examples of $C^{1+\alpha}$ IFSs was chosen precisely to verify the Measure Non-Linearity condition. In light of recent progress on exponential mixing for 33D flows [44] and Axiom $A$ maps [26], it is natural to ask what axiomatic properties of $\tau$ are sufficient to guarantee Quantitative Non-Linearity. More concretely, let $F_{\mu}(x):=\mu(-\infty,x)$ denote the Cantor staircase associated to $\mu$. Since $(F_{\mu})_{*}\mu$ is Lebesgue when $\mu$ is non-atomic, one expects MNL to hold at least when $\tau$ behaves like a Cantor staircase on $F_{\Phi}$; this motivates the example in Theorem 1.4. Generally, MNL quantifies the spread of the values $X_{\mathbf{a}}(\cdot,x_{0})-X_{\mathbf{b}}(\cdot,x_{0})$ on $F_{\Phi}$. Under (MNL) one can show the following variant of UNI: for all $N\geq 1$, there exist $\widehat{\mathbf{a}},\widehat{\mathbf{b}}\in\mathcal{A}^{N}$ such that for all $\mathbf{c}\in\mathcal{A}^{*}$ and all $y\in I_{\mathbf{c}}$ there exists $x\in I_{\mathbf{c}}$ with
    $|X_{\widehat{\mathbf{a}}}(x,y)-X_{\widehat{\mathbf{b}}}(x,y)|\ \geq\ \frac{1}{3C_{\text{(MNL)}}}\,\mu(I_{\mathbf{c}})\qquad\text{(MNL–UNI)}.$
    This is slightly weaker than $((3C_{\text{(MNL)}})^{-1},\delta)$–UNI, and better adapted to truly $C^{1+}$ settings. The non-emptiness of (MNL) suggests prospects for a $C^{1+}$ theory of power Fourier decay and Fractal Uncertainty Principles in the spirit of [13]. It is natural to seek Fourier decay for $C^{1+}$ IFS under UNI-type conditions inspired by (MNL–UNI), and to explore $\tau$ with autosimilar oscillations (e.g. Brownian/Weierstrass-type) which could be used as basis for an axiomatic definition needed for QNL.
  • (2)
    Spectral gaps and Fourier integral operators in low regularity. The motivation for proving Fourier decay is closely tied to its spectral implications, including scattering resonances [23, 13]. Since Theorem 1.1 yields Fourier decay bounds for general phases and slowly growing amplitudes, it is interesting to explore how these methods interact with the spectral theory of transfer and Fourier integral operators in low regularity. In particular, combining our results with ideas from [29] could pave the way to establishing spectral gap properties for twisted transfer operators associated to the $C^{1+\alpha}$ IFSs constructed in Theorem 1.4.
  • (3)
    On quantitative aspects. The decay exponent $\beta$ can be traced explicitly to the exponent in the discretized sum–product theorem 3.2. This depends on a non-concentration exponent that can be written in terms of the (QNL) constants, which in turn can be expressed via the (MNL) or (UNI) constants and the light-tail exponent $\gamma_{0}$. Developing the multiscale induction in the proof of (QNL) under (MNL) or (UNI) is an interesting problem to improve the Fourier decay rates.
  • (4)
    Overlapping IFSs. Our current approach relies on symbolic coding under a separation condition. However, previous work has shown that Fourier decay can also be obtained in overlapping settings via a random Dolgopyat operator method and co-cycle versions of Dolgopyat’s argument [10, 5]. It seems plausible that a variant of the induction-on-scales scheme here (Section 6) could be adapted to handle such overlaps, providing a complementary path to Fourier decay beyond separation.
  • (5)
    Theorem 1.3 is stated for finitely many branches. Extending it to countably many branches (to directly address parabolic examples) is an interesting problem, and would likely require adapting the proof of Theorem 1.2 to infinite alphabets.
1.3. Results in parabolic dynamics
Since Theorem 1.1 applies whenever (QNL) holds, and (QNL) follows from (UNI) in the $C^{2}$ category by Theorem 1.2, our tree strategy avoids using spectral gaps of twisted transfer operators (as in [41, 27, 10, 5]) to obtain power Fourier decay. This framework therefore extends many known results to parabolic systems without having to establish spectral gap estimates. The interface we establish between Fourier–decay estimates and the inducing scheme appears to be new, and it may have further applications in parabolic dynamics in relation to spectral theory; it also treats phases and amplitudes that exhibit slow frequency growth.
In Section 1.3.1, using the coding of Li–Pan [31], we obtain Fourier dimension bounds for Patterson–Sullivan measures on geometrically finite hyperbolic surfaces with cusps. In Section 1.3.2, we treat more general parabolic situations modeled by Young towers, focusing on conformal measures for Manneville–Pommeau maps; the same method applies to other induced systems under analogous distortion and non-integrability assumptions.
1.3.1. Patterson–Sullivan measures on parabolic limit sets
@@
This follows from Li and Pan [31, Prop. 4.1] together with Theorem 1.1. Indeed Li-Pan construct a symbolic coding in which $\mu_{\Gamma}$ is a Gibbs measure with a light-tail property for a subshift of finite type of an IFS $\Phi$ satisfying (1)–(3) in our setting, and $\Lambda_{\Gamma}$ coincides with the attractor $F_{\Phi}$. More precisely, there exists a fundamental interval $I\subset\dot{\mathbb{R}}$ for the parabolic fixed point at $\infty$ (denoted $\Delta_{\infty}$ in [31, §2.3]), constants $C_{1}>0$, $\kappa_{+}\in(0,1)$, $\gamma_{0}>0$, a countable family of disjoint open sets $\{I_{a}\}_{a\in\mathcal{A}}\subset I$, and an expanding map $T:\bigcup_{a}I_{a}\to I$ such that:
  • (1)
    $\sum_{a\in\mathcal{A}}\mu_{\Gamma}(I_{a})=\mu_{\Gamma}(I)$;
  • (2)
  • (3)
    (uniform contraction) $|g_{a}^{\prime}(x)|\leq\kappa_{+}$ for $x\in I$;
  • (4)
  • (5)
With $\phi(x):=-\delta\log|\gamma_{a}^{\prime}(x)|_{\mathbb{S}^{1}}$ for $x\in I_{a}$, equivariance yields that $\mu_{\Gamma}$ is a Gibbs measure for $\phi$, and the $(C_{\phi,\gamma_{0}},\gamma_{0})$–tail condition holds. Finally, Li–Pan prove a UNI statement [31, Lemma 4.5]: writing $\Phi^{n}=\{\gamma_{\mathbf{a}}:\mathbf{a}\in\mathcal{A}^{n}\}$ and $\tau_{n}(x)=\sum_{k=0}^{n-1}\tau(T^{k}x)$, there exist $r,c_{0}>0$ such that for all large $N$ there are words $\{\gamma_{mj}\}_{m=1,2;\,1\leq j\leq j_{0}}$ with
which implies $(c_{0},1)$–UNI by the mean value theorem. Therefore Theorem 1.1 applies, proving Theorem 1.5. The decay exponent in Theorem 1.5 can be traced explicitly in terms of the contraction/distortion data, the UNI constants, and the tail exponent (all available from [31]).
Remark (FUP and resonances). Bourgain–Dyatlov [13] used Fourier decay to prove a Fractal Uncertainty Principle (FUP) for convex co-compact groups ($\delta_{\Gamma}<1/2$), with consequences for the distribution of scattering resonances. In the parabolic case one has $\delta_{\Gamma}\geq 1/2$, and the strategy in [13] does not apply directly. It would be interesting to investigate whether Theorem 1.5 has implications for the meromorphic continuation of the resolvent on geometrically finite manifolds with cusps (cf. [24]).
1.3.2. Parabolic non-linear systems
@@
    Lorenz maps. Consider the interval map $f:J\setminus\{0\}\to J$, J=[1/2,1/2]J=[-1/2,1/2], defined by [6, §3]
@@
    arising from the Poincaré map of the geometric Lorenz flow ($\lambda_{2}<\lambda_{3}<0<-\lambda_{3}<\lambda_{1}$). The derivative of $f$ blows up near 0, and inverse branches have zero derivative there. Diaz-Ordaz [19, Thm. 2] constructs an induced expanding Markov map $T_{A}:I\to I$ (for some neighborhood $I\subset[-1,1]$ of 0) with bounded distortion and first return time $R:I\to\mathbb{N}$ having exponential tails, under uniform contraction (1) and bounded distortion (2).
Induced potentials and summable variations. Given a base potential $\phi_{0}$ for the original system (e.g. $\phi_{0}=-\delta\log|T^{\prime}_{\alpha}|$ for Manneville–Pommeau or $\phi_{0}=-\delta\log|f^{\prime}|$ for Lorenz), define the induced potential on $I$ by $\phi(x):=S_{R(x)}\phi_{0}(x)$. For our results we require summable variations for $\phi$:
$\sum_{n\geq 1}V_{n}(\phi)<\infty,\qquad V_{n}(\phi):=\sup_{\mathbf{a}\in\Sigma^{n}}\ \sup_{x,y\in I_{\mathbf{a}}}|\phi(x)-\phi(y)|,$
which ensures that the Sarig theory applies to the countable Markov map $T_{A}$ [43, 42]. A sufficient condition is $\sum_{n}n\,V_{n}(\phi_{0})<\infty$ [15, §2.4].
Pushing down Gibbs measures. Let $\mu_{\phi}$ be a Gibbs measure for the induced map $T_{A}$. Define the projected measure on the original system by
@@
where $w_{a}(x)=e^{\phi(g_{a}(x))}$. Let $\mu_{\phi}$ be the Gibbs measure for $T_{A}$ arising from the Manneville–Pommeau map $T_{\alpha}$. Then the projected measure $\nu_{\phi}$ for $T_{\alpha}$ satisfies, for some $\beta>0$,
@@
Heavy-tailed measures. Finally, an intriguing direction is to relax the light-tail assumption. Because Theorem 1.1 is formulated in terms of exponential moment conditions, it is well-suited to being adapted to heavier tails, where one might expect sub-polynomial (e.g. polylogarithmic) Fourier decay. In the context of Furstenberg measures, such log–log decay has already been demonstrated in [20]. Extending our framework to this regime would provide a natural bridge between strong polynomial decay and much weaker forms of Fourier decay, potentially revealing new structures in dynamical systems with weaker hyperbolicity.
@@
The paper is dedicated to the proof of Theorem 1.1, Theorem 1.2, Theorem 1.3, Theorem 1.4 and Theorem 1.6. In Section 2 we first give the preliminary thermodynamic notations and derive the needed moment bounds and upper regularity using the light tail condition where the uniform contraction, bounded distortions and light tail conditions are crucially used. In Section 3 we show how the moments arise from the discretised sum-product bounds if we allow the multiplicative random variables have unbounded support. In Section 4 we then begin the proof of Theorem 1.1 by reducing the Fourier transform of the equilibrium state $\mu$ to exponential sums of multiplicative random variables with moment bounds. To then apply the moment bounds and upper regularity derived in Section 2 and the sum-product bound from Section 3, in Section 5 we prove the needed non-concentration bound from (QNL). In section 6, we prove that (QNL) holds under (MNL) and under $(c_{0},\alpha)$-(UNI), which is Theorem 1.2 and Theorem 1.3. In section 7 we construct a $C^{1+}$ IFS satisfying (MNL) and prove Theorem 1.4. Finally, in section 8 we prove Theorem 1.6.
@@ 2. Moment bounds and upper regularity
In this section we introduce some supplementary notations and prove two preliminary statements that will prove useful to prove Theorem 1.1. The first is a precise bound coming from the properties of the pressure. The second is an upper-regularity estimate for $\mu_{\phi}$ on this setting.
@@ 2.1. Symbolic coding, Gibbs measures and transfer operators
Recall the notation from Section 1.2, where we are given open disjoint intervals $I_{a}\subset[0,1]$, $a\in\mathcal{A}$ and a $C^{1}$ map $T:\overline{I}\to\mathbb{R}$ for $I=\bigcup_{a\in\mathcal{A}}I_{a}$ satisfying the Markov property: for all $a,b$ such that we have that $T(\overline{I_{b}})\cap(I_{a})\neq\emptyset$, then $T(\overline{I_{b}})\supset\overline{I_{a}}$, which gives us a $\{0,1\}$ transition matrix $A$ satisfying $A(a,b)=1$ if and only if $T(\overline{I_{a}})\supset\overline{I_{b}}$. This gives us a coding $\Sigma^{\mathbb{N}}:=\{\mathbf{a}\in\mathcal{A}^{\mathbb{N}}:A(a_{j},a_{j+1})=1\}$, where length $n$ words are denoted by $\Sigma^{n}$ and finite words are denoted $\Sigma^{*}=\bigcup_{n\in\mathbb{N}}\Sigma^{n}$. Then a word $\mathbf{a}:=a_{1}\dots a_{n}\in\mathcal{A}^{n}$ satisfying $A(a_{i},a_{i+1})=1$ for all $i$ is called admissible and for each admissible $ab\in\mathcal{A}^{2}$, we have an associated inverse branch $g_{ab}:I_{b}\to I_{a}$, which form an iterated function system $\Phi=\{g_{ab}:ab\in\mathcal{A}^{2}\ \text{admissible}\}$. Recall that the notation $\mathbf{a}\to x$ simply means that $g_{\mathbf{a}}$ is defined at $x$, i.e. $x\in I_{b\left(\mathbf{a}\right)}$ with $b\left(\mathbf{a}\right)$ denoting the last letter of $\mathbf{a}$.
We then assume the IFS $\Phi=\{g_{ab}:ab\in\mathcal{A}^{2}\ \text{admissible}\}$ satisfies the conditions (1)-(3) from Section 1.2, though in this section we only need uniform $\kappa_{+}$ contraction (1) and bounded $(C,\alpha)$-distortions (2). The uniform contraction of the maps $g_{ab}$ guarantees that there exists a unique attractor $F\subset I$ associated to $\Phi$, that is, for $\mathbf{a}\in\Sigma^{n+1}=a_{0}\dots a_{n}$ the map $g_{\mathbf{a}}:=g_{a_{0}a_{1}}\circ\dots\circ g_{a_{n-1}a_{n}}:I_{a_{n}}\rightarrow I_{a_{0}}$, and the interval $I_{\mathbf{a}}:=g_{\mathbf{a}}(I_{a_{n}})$, then $F$ is the compact set given by $F=\bigcap_{n\in\mathbb{N}}\bigcup_{\mathbf{a}\in\Sigma^{n}}\overline{I_{\mathbf{a}}}$.
Let $\mu=\mu_{\phi}$ be the Gibbs measure given by some normalized potential $\phi$, which is uniquely defined to satisfy the $\phi$-conformality condition:
@@
for all continuous $f:F\to\mathbb{C}$, where $b(\mathbf{a})$ is the last letter of a word, so that we can write $g_{\mathbf{a}}:I_{b(\mathbf{a})}\rightarrow I_{\mathbf{a}}$. We also assume $\mu_{\phi}$ satisfies the light $(C_{\phi,\gamma_{0}},\gamma_{0})$-tail condition (4). Now the $\phi$-conformality condition can be iterated to yields the formula:
@@
where $S_{n}\phi(x):=\sum_{k=0}^{n-1}\phi(T^{k}(x))$ is a Birkhoff sum. It is then convenient to denote $w_{\mathbf{a}}(x):=e^{S_{n}\phi(g_{\mathbf{a}}(x))}$. The invariance formula can then be rewritten as
@@
$\forall\mathbf{a}\in\Sigma^{n},\forall x\in I_{b(\mathbf{a})},\ C_{\phi}^{-1}w_{\mathbf{a}}(x)\leq p_{\mathbf{a}}:=\mu(I_{\mathbf{a}})\leq C_{\phi}w_{\mathbf{a}}(x).$
Troughout the text well will use the notation $a_{n}\lesssim b_{n}$ if there exists some constant $C\geq 1$ such that $a_{n}\leq Cb_{n}$. We will write $a_{n}\simeq b_{n}$ if $a_{n}\lesssim b_{n}\lesssim a_{n}$. Gibbs estimates are then rewritten $w_{\mathbf{a}}(x)\simeq\mu(I_{\mathbf{a}})=p_{\mathbf{a}}$ for all $x$.

### v3

@@ 1.2. QNL and power Fourier decay
For all $0<t<1$, the Lyons conductance measure $\nu_{t}$ has power Fourier decay.
Now Shmerkin’s strategy [70] cannot be directly used with Theorem 1.3 to prove Lyons’s conjecture due to the lack of convolution structure of $\nu_{t}$. However, as in the work of Lindenstrauss and Varjú on the higher dimensional setting [50], which is very different from the one dimensional case due to non-Abelian nature of $SO(d)$ instead a spectral gap assumption on the underlying transfer operator was crucial in the proof of absolute continuity via a power Fourier decay property of the self-similar measure. Given that our proof gives a non-concentration for the Lyons IFS typically needed for spectral gaps [13], one may try to adapt more directly the spectral gaps of twisted transfer operators associated to $\nu_{t}$ as in [50] to try to upgrade Theorem 1.3 to a proof of Lyons conjecture.
1.4. Patterson–Sullivan measures on parabolic limit sets.
Fourier decay for Patterson–Sullivan measures associated to convex co-compact hyperbolic surfaces was established by Bourgain and Dyatlov [18] as a tool to bound the resonance free regions for scattering resonances by proving improved the Fractal Uncertainty Principle bounds. Extension of this to convex co-compact subgroups of $\mathrm{PSL}(2,\mathbb{C})$ was done in [47] using twisted transfer operators and to higher dimensions in [12] adapting an $L^{2}$-flattening approach of Khalil [40]. We will now see that Theorem 1.1 allows us to expand these to geometrically finite case where there are parabolic elements in the group.
@@
Bourgain–Dyatlov [18] used Fourier decay to prove a Fractal Uncertainty Principle (FUP) for convex co-compact groups ($\delta_{\Gamma}<1/2$), with consequences for the distribution of scattering resonances. In the parabolic case one has $\delta_{\Gamma}\geq 1/2$, and the strategy in [18] does not apply directly. It would be interesting to investigate whether Theorem 1.4 has implications for the meromorphic continuation of the resolvent on geometrically finite manifolds with cusps (cf. [36]).
1.5. Manneville–Pommeau and Lorenz maps.
@@
    Lorenz maps. Consider the map $f:J\setminus\{0\}\to J$, J=[1/2,1/2]J=[-1/2,1/2], defined by [8, §3]
@@
    arising from the Poincaré map of the geometric Lorenz flow ($\lambda_{2}<\lambda_{3}<0<-\lambda_{3}<\lambda_{1}$). The derivative of $f$ blows up near 0, and inverse branches have zero derivative there. Diaz-Ordaz [29, Thm. 2] constructs an induced expanding Markov map $T_{A}:I\to I$ (for some neighborhood $I\subset[-1,1]$ of 0) with bounded distortion and first return time $R:I\to\mathbb{N}$ having exponential tails, under uniform contraction (1) and bounded distortion (2).
Given a base potential $\phi_{0}$ for the original system (e.g. $\phi_{0}=-\delta\log|T^{\prime}_{\alpha}|$ for Manneville–Pommeau or $\phi_{0}=-\delta\log|f^{\prime}|$ for Lorenz), define the induced potential on $I$ by $\phi(x):=S_{R(x)}\phi_{0}(x)$. For our results we require summable variations for $\phi$: $\sum_{n\geq 1}V_{n}(\phi)<\infty$, where $V_{n}(\phi):=\sup_{\mathbf{a}\in\Sigma^{n}}\ \sup_{x,y\in I_{\mathbf{a}}}|\phi(x)-\phi(y)|,$ which ensures that the Sarig theory applies to the countable Markov map $T_{A}$ [69, 68]. A sufficient condition is $\sum_{n}n\,V_{n}(\phi_{0})<\infty$ [20, §2.4]. Let $\mu_{\phi}$ be a Gibbs measure for the induced map $T_{A}$. Define the projected measure on the parabolic system by
@@
where $w_{a}(x)=e^{\phi(g_{a}(x))}$. Let $\mu_{\phi}$ be the Gibbs measure for $T_{A}$ arising from the Manneville–Pommeau map $T_{\alpha}$ or Lorenz map $f$. Then the projected parabolic measure $\nu_{\phi}$ for $T_{\alpha}$ (or Lorenz map $f$ respectively) satisfies, for some $\beta>0$,
@@
An intriguing direction is to relax the light-tail assumption. Because Theorem 1.1 is formulated in terms of exponential moment conditions, it is well-suited to being adapted to heavier tails, where one might expect sub-polynomial (e.g. polylogarithmic) Fourier decay. In the context of Furstenberg measures, such log–log decay has already been demonstrated in [30]. Extending our framework to this regime would provide a natural bridge between strong polynomial decay and much weaker forms of Fourier decay, potentially revealing new structures in dynamical systems with weaker hyperbolicity.
1.6. $C^{1+\alpha}$ iterated function systems.
Finally, many dynamically defined measures of interest arise from systems that admit only $C^{1+\alpha}$ regularity, for instance due to rough holonomies in codings of hyperbolic diffeomorphisms or in nonlinear cocycle models such as circle extensions and the Fibonacci Hamiltonian. In such settings the associated temporal distance functions are typically non-smooth, and the classical Fourier decay mechanisms based on Dolgopyat-type spectral gap estimates are not available.
Our framework applies in this genuinely low-regularity regime. In particular, we construct examples of $C^{1+\alpha}$ iterated function systems whose conformal measures admit power Fourier decay despite not being $C^{1}$-conjugate to linear models, that is, there is no $C^{1}$ map $h$ such that $\{hfh^{-1}:f\in\Phi\}$ consists of maps with constant derivative.
Theorem 1.6.
There exist $C^{1+\alpha}$ iterated function systems with finitely many branches whose attractors support equilibrium states with power Fourier decay, and which are not $C^{1}$-conjugate to linear IFS.
In particular, these examples provide conformal measures of positive Fourier dimension arising from genuinely $C^{1+\alpha}$ dynamics. The construction proceeds by coupling the symbolic and geometric oscillations of the distortion function through a novel fixed-point argument in the moduli space of IFS parametrised by equilibrium states. This allows us to verify the multiscale nonlinearity condition required by Theorem 1.1 without relying on higher regularity or smooth symbolic holonomies.
The condition on MNL is quite restrictive as it requires oscillations at right dimension on the fractal. This suggests the possibility of developing a $C^{1+}$ theory of Fourier decay, spectral gaps and fractal uncertainty principles in settings where the distortion functions exhibit only autosimilar or fractal-type oscillations (e.g. Brownian/Weierstrass-type as in [76, 42]), rather than smooth curvature.
@@
In Section 2 we first give the preliminary thermodynamic notations and derive the needed moment bounds and upper regularity using the light tail condition where the uniform contraction, bounded distortions and light tail conditions are crucially used. In Section 3 we show how the moments arise from the discretised sum-product bounds if we allow the multiplicative random variables have unbounded support. In Section 4 we then begin the proof of Theorem 1.1 by reducing the Fourier transform of the equilibrium state $\mu$ to exponential sums of multiplicative random variables with moment bounds. To then apply the moment bounds and upper regularity derived in Section 2 and the sum-product bound from Section 3, in Section 5 we prove the needed non-concentration bound from (QNL). In section 6, we prove that (QNL) holds under UNI and MNL conditions. In Section 7 we check conditions for Theorem 1.3. In section 8 we prove Theorem 1.4. In section 9 we prove Theorem 1.5. In section 10 we construct a $C^{1+}$ IFS satisfying (MNL) and prove Theorem 1.6.
@@ 2. Moment bounds and upper regularity
In this section we introduce some supplementary notations and prove two preliminary statements that will prove useful to prove Theorem 1.1. The first is a precise bound coming from the properties of the pressure. The second is an upper-regularity estimate for $\mu_{\phi}$ on this setting. A key reference to the thermodynamical formalism for countable Markov shift we use in the symbolic level is given by Sarig’s papers [69, 68].
@@ 2.1. Symbolic coding, Gibbs measures and transfer operators
Recall the notation from Section 1.2, where $\Sigma^{\mathbb{N}}:=\{(a_{1},a_{2},\ldots)\in\mathcal{A}^{\mathbb{N}}:\ A(a_{i},a_{i+1})=1\ \text{for all }i\}$ is the countable Markov shift determined by a matrix $A\in\{0,1\}^{\mathcal{A}\times\mathcal{A}}$ and $\Sigma^{n}$ is the set of admissible words of length $n$ and by $\Sigma^{*}=\bigcup_{n\geq 1}\Sigma^{n}$. Then a word $\mathbf{a}:=a_{1}\dots a_{n}\in\mathcal{A}^{n}$ satisfying $A(a_{i},a_{i+1})=1$ for all $i$ is called admissible and cylinder $\mathbf{a}]\subset\Sigma^{\mathbb{N}}$ is the subtree of infinite words with initial $\mathbf{a}$. We assume there is a collection $\{I_{a}\}_{a\in\mathcal{A}}$ of possibly overlapping open intervals in $[0,1]$ and for each admissible pair $ab\in\Sigma^{2}$ we are given a $C^{1+\alpha}$ contraction $g_{ab}:I_{b}\to I_{a},$ which form an iterated function system $\Phi=\{g_{ab}:ab\in\mathcal{A}^{2}\ \text{admissible}\}$. For an admissible word $\mathbf{a}=a_{1}\cdots a_{n}\in\Sigma^{n}$ we define $g_{\mathbf{a}}:=g_{a_{1}a_{2}}\circ\cdots\circ g_{a_{n-1}a_{n}}.$ Recall that the notation $\mathbf{a}\to x$ simply means that $g_{\mathbf{a}}$ is defined at $x$, i.e. $x\in I_{b\left(\mathbf{a}\right)}$ with $b\left(\mathbf{a}\right)$ denoting the last letter of $\mathbf{a}$.
We suppose the IFS $\Phi=\{g_{ab}:ab\in\mathcal{A}^{2}\ \text{admissible}\}$ satisfies the conditions (1)-(3) from Section 1.2, though in this section we only need uniform $\kappa_{+}$ contraction (1) and bounded $(C,\alpha)$-distortions (2). The uniform contraction of the maps $g_{ab}$ guarantees that there exists a unique attractor $F\subset I$ associated to $\Phi$, that is, for $\mathbf{a}\in\Sigma^{n+1}=a_{0}\dots a_{n}$ the map $g_{\mathbf{a}}:=g_{a_{0}a_{1}}\circ\dots\circ g_{a_{n-1}a_{n}}:I_{a_{n}}\rightarrow I_{a_{0}}$, and the interval $I_{\mathbf{a}}:=g_{\mathbf{a}}(I_{a_{n}})$, then $F$ is the compact set given by $F=\bigcap_{n\in\mathbb{N}}\bigcup_{\mathbf{a}\in\Sigma^{n}}\overline{I_{\mathbf{a}}}$. This also gives us a projection $\pi:\Sigma^{\mathbb{N}}\to F$.
Since the IFS $\Phi$ we work on in Theorem 1.1 has overlaps, the map $\pi$ can be very badly away from being injective, we will need to consider thermodynamical formalism and the associated measures $\mu$ as projections of measures on the symbolic space $\Sigma^{\mathbb{N}}$. Formally, if $\phi:F\to\mathbb{R}$ is a potential, we can consider $\phi$ also as a map $\phi\circ\pi:\Sigma^{\mathbb{N}}\to\mathbb{R}$. Then if we let $\nu_{\phi}$ be the (symbolic) Gibbs measure associated to $\phi$ on $\Sigma^{\mathbb{N}}$, we mean that if is a probability measure satisfying
$\int_{\Sigma^{\mathbb{N}}}f(\mathbf{a})\,d\nu_{\phi}(\mathbf{a})=\sum_{ab\in\Sigma^{2}}\int_{[b]}e^{\phi(g_{ab}(\mathbf{a}))}f(g_{ab}(\mathbf{a}))\,d\nu_{\phi}(\mathbf{a})$
for all continuous $f:\Sigma^{\mathbb{N}}\to\mathbb{C}$. Here $b(\mathbf{a})$ is the last letter of a word. Thus this implies the projected Gibbs measure $\mu_{\phi}:=\pi\mu_{\phi}$ on the attractor $F$ satisfies the $\phi$-conformality condition
@@
for all continuous $f:F\to\mathbb{C}$ and note by the definition of $b(\mathbf{a})$ we can write $g_{\mathbf{a}}:I_{b(\mathbf{a})}\rightarrow I_{\mathbf{a}}$. So from now on when we write $\mu_{\phi}$ is a Gibbs measure, we assume $\mu_{\phi}$ comes as a projection of a symbolic measure $\nu_{\phi}$ defined on $\Sigma^{\mathbb{N}}$. We also assume $\mu_{\phi}$ satisfies the light $(C_{\phi,\gamma_{0}},\gamma_{0})$-tail condition (4) from Section 1.2.
Now the $\phi$-conformality condition can be iterated to yields the formula:
@@
where $S_{n}\phi(g_{\mathbf{a}}(x)):=\sum_{k=0}^{n-1}\phi(g_{\mathbf{a}|k}(x))$ is a Birkhoff sum. It is then convenient to denote $w_{\mathbf{a}}(x):=e^{S_{n}\phi(g_{\mathbf{a}}(x))}$. The invariance formula can then be rewritten as
@@
$\forall\mathbf{a}\in\Sigma^{n},\forall x\in I_{b(\mathbf{a})},\ C_{\phi}^{-1}w_{\mathbf{a}}(x)\leq p_{\mathbf{a}}:=w_{\mathbf{a}}(x_{b(\mathbf{a})})\leq C_{\phi}w_{\mathbf{a}}(x).$
Troughout the text well will use the notation $a_{n}\lesssim b_{n}$ if there exists some constant $C\geq 1$ such that $a_{n}\leq Cb_{n}$. We will write $a_{n}\simeq b_{n}$ if $a_{n}\lesssim b_{n}\lesssim a_{n}$. Gibbs estimates are then rewritten $w_{\mathbf{a}}(x)\simeq p_{\mathbf{a}}$ for all $x$.